OpenStudy (anonymous):

solve 4x-4y=-13 and 2x-4=-2

OpenStudy (anonymous):

y=-4/9. x=1 if your question is correct, which i think it is not!

OpenStudy (anonymous):

x=1, y=17/4

OpenStudy (anonymous):

is this eqn right!!!!!

OpenStudy (anonymous):

whats wrong with this first equation is an inclined line the 2nd is a vertical one

jhonyy9 (jhonyy9):

if this is write correct ,so than 4x -4y = -13 2x-4 = -2 --- 2x= -2 +4 --- 2x=2 --- x=2/2 --- x=1 if x=1 in the first equation we will get 4 -4y =-13 --- 17= 4y --- y= 17/4